find the coordinates of the midpoint of each side.


A(3,6) B(-20) C(3,1)
midpoint of AC?
midpoint of AB?
midpoint of BC?

Answer by checkley71(8403)

A(3,6) B(-2,0)? C(3,1)
6+0/2=3 6-3=3 FOR THE Y VALUE, 2+3=5/2=2.5 3-2.5=.5 FOR THE X VALUE OR (.5,3)
1=0/2=.5 1-.5=.5 FOR THEY VALUE, 3+2=5/2=2.5 3-2.5=,5 FOR THE X VALUE OR (.5,.5)
6-1=5/2=2.5 6-2.5=3.5 FOR THE Y VALUE, 3-3=0 3-0=3 FOR THE X VALUE OR (3,3.5)
MIDPOINT AB=(.5,3)
MIDPOINT BC=(.5,.5)
MIDPOINTAC=(3,3.5)